40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 31 20:06:22 2013, Evan, Update, Locking, PRM/PR2 cavity 
    Reply  Thu Jan 31 23:48:45 2013, Koji, Update, Locking, PRM/PR2 cavity 
       Reply  Fri Feb 1 09:33:11 2013, Jamie, Update, Locking, PRM/PR2 cavity 
          Reply  Fri Feb 1 15:12:53 2013, Jenne, Update, Locking, PRM/PR2 cavity 
             Reply  Fri Feb 1 19:55:33 2013, Jenne, Update, Locking, PRM/PR2 cavity 
                Reply  Fri Feb 1 23:12:42 2013, Koji, Update, Locking, PRM/PR2 cavity 
                Reply  Fri Feb 1 23:52:59 2013, rana, Update, Locking, PRM/PR2 cavity 
Message ID: 7981     Entry time: Fri Feb 1 09:33:11 2013     In reply to: 7980     Reply to this: 7985
Author: Jamie 
Type: Update 
Category: Locking 
Subject: PRM/PR2 cavity 

I replaced the BS1 between the POPDC PD and the camera with a 98 reflector, and moved the 50 up before the BS to dump half the light.  Still saturating POPDC, but hopefully the ratio between POPDC and the camera should be better.  We just need to dump more of the power before we get there.  I'll come back to this after C&D if no one else has already gotten to it.

I don't know why I didn't pay more attention last night, but things look way WAY better.  The beams are much cleaner and the power level is much much higher.

ELOG V3.1.3-